OpenStudy (anonymous):

show that f is the function defined by f(x) = mx+b where m is not equal to 0, then the inverse function f^-1 is defined by the formula f^-1(y) =(1/m)y - (b/m)

OpenStudy (anonymous):

solve \[y=mx+b\] for \(x\)

OpenStudy (anonymous):

y-b/m = x

OpenStudy (anonymous):

right? @satellite73

OpenStudy (anonymous):

then (y-b)/m = (1/m)y-(b/m)

OpenStudy (anonymous):

then (y-b)/m = (y/m)-(b/m) which is (y-b)/m

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I don't buy it. You haven't shown that the range of f(x) is the domain of f^-1 or the domain of f(x) is the range of f^-1(y) In other words, you need to show that f(f^-1(y)) = y and that f^-1(f(x)) =x